Circular No. 5770

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993J IN NGC 3031 V. Klochkova and V. Panchuk, Special Astrophysical Observatory, re-

port: “Four CCD spectrograms (range 400-1000 nm; resolution ∼ 0.15-0.2 nm, S/N > 100) of SN 1993J were obtained on Apr. 5/6 at the 6-m telescope (+ medium-resolution échelle spectrometer, Zebra). The Balmer and Pashen series of hydrogen are absent in these spectra. The only spectral details revealed were the interstellar Na I 588.9- and 589.5-nm lines, each of these doublet components being split; we estimated the radial velocities for the components as –103, –2, and +128 km/s.”

P. Prugniel, Observatoire de Haute Provence, writes: “To match the ‘zero point’ of Corwin’s photometry for star 0928 (IAUC 5742), corrections of –0.11 in B_{c} and +0.04 in B_{c}–R_{c} should be added to my measurements published since IAUC 5736. The supernova is now fading in B, V, R, and I. The shape of the lightcurve is almost symmetrical around the secondary maxima, which from the OHP monitoring are as follows: B_{c}, 11.51, Apr. 16.7 UT; V_{c}, 10.87, 18.1; R_{c}, 10.52, 19.0; I_{c}, 10.43, 19.6.”

E. Guinan, J. Marshall, and G. McCook, Villanova University, report on UBVRI photoelectric photometry of SN 1993J, conducted with the FCC 0.8-m Automatic Photometry Telescope on Mt. Hopkins. The observations have been obtained since Apr. 2 UT using HD 85458 and HD 86677 as comparison stars. The observations show a local light minimum at Apr. 5.1 ± 0.5. The star has gone through a second wavelength-dependent light maximum. The times of the second light maximum in the U, B, V, R, and I bandpasses are Apr. 15.2, 16.5, 18.0, 18.6, and 19.0, respectively. After Apr. 15, SN 1993J has rapidly decreased in brightness in the U bandpass.

The following photometry (cf. IAUC 5750, 5755, 5758) by M. Okyudo and T. Ishida, Nishi-Harima Astronomical Observatory (OI; generally ± 0.01 mag); by Guinan et al. (G; see above); and by Prugniel (P; see above), are Johnson V magnitudes unless otherwise noted: Apr. 2.12 UT, 11.46 (G); 3.13, 11.67 (G); 4.13, 11.81 (G); 5, 11.84 (G); 7.14, 11.71 (G); 9.14, 11.53 (G); 10.13, 11.32 (G); 10.608, 11.46 (OI); 11.659, 11.30 (OI); 12.557, 11.24 (OI); 13.577, 11.17 (OI); 14.628, 11.03 (OI); 16.14, 11.00 (G); 16.601, 10.88 (OI); 17.15, 10.97 (G); 17.612, 11.00 (OI); 18.644, 10.97 (OI); 19.15, 10.98 (G); 19.634, 10.85 (OI); 20.16, 11.01 (G); 20.605, 11.04 (OI); 21.00, V_{c} = 10.96, B_{c}–V_{c} = +0.76, V_{c}–R_{c} = +0.40, V_{c}–I_{c} = +0.50 (P).

1993 April 21 Daniel W. E. Green